Using the formula: h = vt - 16t^2 , find the time that a rocket reaches a height of 144 ft if its v = 96 ft/sec
Answer by rothauserc(4718)

h = vt - 16t^2
144 = 96t - 16t^2
16t^2 -96t +144 = 0
divide both sides of = by 16
t^2 -6t +9 = 0
(t -3) = 0
t = 3 seconds